OpenStudy (amenah8):

How would you solve for p in this inequality? 7≥9–p

jagr2713 (jagr2713):

Hey what do you think we do first?

OpenStudy (amenah8):

First, like any equation, we would subtract 9 from both sides, right?

jagr2713 (jagr2713):

Coreect so what do we get?

OpenStudy (kash_thesmartguy):

We take the constant off. So subtract 9 from both sides. There is your answer

jagr2713 (jagr2713):

No @Kash_TheSmartGuy

OpenStudy (amenah8):

-2 (greater than/equal to sign) -p would be the answer?

OpenStudy (kash_thesmartguy):

No, - and - equals positive. your answer will be in positive

jagr2713 (jagr2713):

so we subtract and get -2 so now we have'' −2≥−p now Multiply both sides by −1 what do we get?

OpenStudy (kash_thesmartguy):

2 is greater than or equal to p

OpenStudy (amenah8):

I agree,the answer is 2 is greater than or equal to p, but IXL says that p is greater than or equal to 2. WHy are they doing this? IXL is adding p to both sides.

OpenStudy (kash_thesmartguy):

IXL is wrong!

jagr2713 (jagr2713):

Yes i agree to its 2 :D Its just that she said subtract 9 and thats your answer but you forgot to multiply by -1

OpenStudy (amenah8):

Yeah, it's wrong on MANY occasions. Thanks, everyone!

OpenStudy (kash_thesmartguy):

np :)

OpenStudy (anonymous):

\[-2 \ge-p\] multiply by -1 \[2\le p\] or \[p \ge 2\]